# Towards more realistic evaluation of LLM-based code generation: an experimental study and beyond

### v1

@@ Abstract.
To evaluate the code generation capabilities of Large Language Models (LLMs) in complex real-world software development scenarios, many evaluation approaches have been developed. They typically leverage contextual code from the latest version of a project to facilitate LLMs in accurately generating the desired function. However, such evaluation approaches fail to consider the dynamic evolution of software projects over time, which we refer to as evolving-ignored situation, leading to issues of future context leakage and useful context missing. This in turn results in inaccurate evaluation of LLMs’ performance. In this paper, we conduct an empirical study to deeply understand LLMs’ code generation performance within settings that reflect the evolving nature of software development. To achieve this, we first construct an evolving-aware repository-level code generation dataset, namely HumanEvo, equipped with an automated execution-based evaluation tool. Second, we manually categorize HumanEvo according to dependency levels to more comprehensively analyze the model’s performance in generating functions with different dependency levels. Third, we conduct extensive experiments on HumanEvo with seven representative and diverse LLMs to verify the effectiveness of the proposed benchmark. We obtain many important findings through our experimental study. For example, we find that previous evolving-ignored evaluation approaches lead to inflated performance of the LLMs, ranging from 10.0% to 61.1%. Based on the findings, we give actionable suggestions on more realistic evaluation of LLMs on code generation. We also build a shared evolving-aware code generation toolbox to facilitate future research. Replication package including source code, datasets and appendix is available at https://github.com/DeepSoftwareAnalytics/EvoEval.
conference: ACM SIGSOFT International Symposium on Software Testing and Analysis; 16-20 September, 2024; Vienna, Austria
1. Introduction
In recent years, the task of code generation based on LLMs has drawn widespread attention (Zheng et al., 2023b, a; Brown et al., 2020; Dong et al., 2023; Jiang et al., 2023; Zhang et al., 2023a; Hong et al., 2024; Qian et al., 2023). Many code LLMs (Rozière et al., 2024; Li et al., 2023; Guo et al., 2024; Zheng et al., 2023c; Chowdhery et al., 2023; Allal et al., 2023; Luo et al., 2024; Christopoulou et al., 2022; Shen et al., 2023; Nijkamp et al., 2023; Chen et al., 2021b; Li et al., 2022; Fried et al., 2023) are being employed to empower programming assistants, which are playing a vital role in practical software development.
@@ 1. Introduction
Recently, several repository-level code generation benchmarks have emerged to better simulate real-world development scenarios, such as CoderEval (Yu et al., 2024), RepoBench (Liu et al., 2024), EvoCodeBench (Li et al., 2024a), and RepoEval (Zhang et al., 2023b). They sample functions from real-world projects as programming questions, aiming to reflect the performance of LLMs in actual development by prompting them to generate these functions according to the target function description and project context.
Figure 1. Issues of future context leakage and useful context missing
However, current benchmarks ignore the evolution nature (Voinea et al., 2005; Cleland, 2004; Robles et al., 2005) of real-world projects by treating the latest version of the repository as context, leading to inaccurate reflection of real development scenario. We identify two issues caused by this inaccurate reflection as depicted in Figure  1. The upper part of Figure 1 shows that, over time, the project repository has undergone multiple commits, including adding new functions and removing functions. In the lower part of Figure 1, we showcase the two issues that may arise from ignoring the evolution nature of a project, which we refer to as evolving-ignored situation. Example 1 illustrates the future context leakage issue that the evolving-ignored situation may lead to. Specifically, in the latest repository version, a code snippet committed in 2019 is highly similar to another snippet committed in 2022. Suppose the code snippet from 2019 is selected as the target function for LLMs. In that case, the code snippet from 2022 should not be included as project context during context acquisition even though they share high similarity. This is because the developer did not have access to the code committed in 2022 when writing the target function. Therefore, providing this code snippet as context to the model could pose the issue of future context leakage, resulting in an overestimated evaluation of LLMs’ code generation performance. Example 2 illustrates the useful context missing issue that the evolving-ignored situation may lead to. The target function is highly similar to the function in the pink box, and programmer deletes the function in the pink box after committing the target function. This suggests that when developers write the added code snippet, they can refer to that similar code snippet for inspiration. However, if the latest version of the project is utilized as context source, the model may fail to generate the target function correctly due to the absence of this useful context. This leads to an underestimation of LLMs’ code generation performance.
@@
In this paper, we conduct an empirical study to comprehensively reveal LLMs’ repository-level code generation performance. To accomplish this goal, we construct HumanEvo, a novel evolving-aware repository-level code generation benchmark, which better simulates real-world development processes and reflects the evolving nature of projects over time. Specifically, projects evolve continuously throughout the development process, meaning that the project context available to programmers when writing different functions is constantly changing. To ensure that the context we provide precisely encompasses all the context available to the programmer when completing the target function (neither including additional code introduced by subsequent version changes nor omitting code that may have been updated or deleted due to version change), we roll back the entire project repository to the state before the target code was committed and evaluate the performance of LLMs on the rolled-back project.
@@
We conduct a rigorous data collection pipeline to construct HumanEvo. We start with selecting high-quality projects and then collecting a large number of pull requests (PRs) from these projects. Then, we perform initial filtering on the crawled PRs to meet three attributes to guarantee the quality of the newly added functions in PRs. Additionally, to ensure that the newly added functions are covered by the test suite, we establish a runtime environment for each PR and run the corresponding project’s testing framework. After all this, we obtain 400 programming questions, 200 for Python and 200 for Java. In our benchmark construction process, our selection of functions revolves around PRs rather than commits. First, functions in a merged PR are typically reviewed by project managers (Rahman and Roy, 2014a), ensuring the quality of the added functions. Second, a PR usually incorporates test code for the added functions, guaranteeing their coverage by the project’s testing framework.
We conduct an extensive evaluation of multiple mainstream LLMs (the CodeLlama series (Rozière et al., 2024), DeepSeekCoder series (Guo et al., 2024), GPT-4 (OpenAI, 2024),GPT-3.5-Turbo) on HumanEvo. Experimental results reveal the performance of LLMs in more practical development scenarios and emphasize the importance of evolving-awareness, showcasing how LLMs gradually deviate from reality in evolving-ignored situations. Specifically, we first compare the performance of LLMs under evolving-ignored and evolving-aware settings, and experimental results indicate that LLMs exhibit inflated performance under the evolving-ignored setting. To investigate whether both issues depicted in Figure 1 have certain impacts on the models, we further analyze task instances successfully completed by the models under both settings and find that both “future context leakage” and “useful context missing” have significant impacts on the models. According to the empirical results, we highlight four key findings as follows:
  1. (1)
    Previous evolving-ignored evaluation approaches lead to inflated performance of the LLMs, ranging from 10.0% to 61.1%, revealing how previous evaluation methods provided the models with a deviated programming scenario from reality.
  2. (2)
    LLMs’ success rates in generating functions with intra-class dependencies decreased by approximately 14.0% on average, and those with inter-class dependencies by approximately 30.9%, under the evolving-aware setting compared to the evolving-ignored setting. This suggests that functions with broader dependencies are more impacted on different evaluation settings.
  3. (3)
    Under fine-granularity evolving-ignored settings, the discrepancy between model performance and real-world effectiveness increases over time, exhibiting a progressive, upward trend in inflated evaluation metrics as projects evolve.
  4. (4)
    Detailed docstrings significantly enhance the LLMs’ ability to accurately generate the expected code.
In summary, our contributions can be summarized as follows:
@@
  • We have identified a common flaw in previous repository-level code generation benchmarks: they are all evolving-ignored, causing inaccurate simulations of real-world development scenarios.
@@
  • We introduce a new benchmark HumanEvo, an evolving-aware repository-level code generation benchmark that addresses the evolving-ignored issue, to conduct a comprehensive empirical study on the performance of LLMs in repository-level code generation.
@@
  • Through extensive experimentation, we substantiate the evolving-ignored situation leads to inflated performance of the models, with LLMs showing a performance inflation ranging from 10.0% to 61.1%.
@@
  • For convenient usage of HumanEvo, we release all data, code, and the evaluation platform based on Docker image that provides runtime environments for all projects, facilitating the automated execution of the test suites.
2. Background
2.1. Code Generation Benchmark
Previously, the vast majority of benchmarks were constructed to evaluate the performance of LLMs in generating standalone functions(Athiwaratkun et al., 2023; Cassano et al., 2022a; Jain et al., 2022; Zan et al., 2022b; Du et al., 2023; Ding et al., 2024; Khan et al., 2023), with the most representative ones being HumanEval  (Chen et al., 2021a) and MBPP  (Austin et al., 2021). HumanEval comprises 164 manually crafted programming problems covering language comprehension, reasoning, algorithms, and simple mathematics. MBPP consists of 974 short Python function problems primarily designed for novice programmers.
Finding previous benchmarks fail to reflect complex programming scenarios, researchers propose CoderEval (Yu et al., 2024), the first repository-level code generation benchmark, to evaluate the performance of existing LLMs in real-world development by prompting them to generate functions selected in actual complex software. CoderEval meticulously selects 230 Python and 230 Java programming tasks from popular real-world open-source projects, evaluating the functional correctness of generated code based on execution. To better simulate real-world development scenarios, EvoCodeBench (Li et al., 2024a) first analyses the distribution of standalone and non-standalone functions in a large number of real-world projects, then constructs a benchmark with a dependency distribution more in line with real-world scenarios.
Although the aforementioned works prompt LLMs to generate functions selected from real software projects to evaluate LLM performance in real development scenarios, we recognize a common flaw in all these benchmarks: they ignore the evolution nature of project repositories. This implies that when we prompt LLMs to generate different functions, the project context provided to LLMs should also consider the evolution of the project. Taking the example of software development processes based on GitHub project management, when writing functions, the project context that programmers can refer to is the entire project repository up to the commit preceding the target functions. Therefore, when prompting LLM to generate target functions, we should roll back the entire project repository to the state before the target function was committed.
2.2. Repository-level Code Generation
Many technical studies have been proposed to better tackle repository-level code generation tasks (Shrivastava et al., 2024; Zhou et al., 2023; Zhang et al., 2023b; Zan et al., 2023, 2022a; Bairi et al., 2023; Liang et al., 2024; Eghbali and Pradel, 2024; Zhang et al., 2024; Phan et al., 2024; Wang et al., 2024). DocPrompting (Zhou et al., 2023) uses natural language (NL) intent to retrieve documents related to the target code and then sends NL intent and retrieved documents to LLMs. CoCoMIC (Ding et al., [n. d.]) develops a cross-file context finder, which is a static code analysis tool used to retrieve the most related cross-file context. MGD (Agrawal et al., 2024) leverages IDEs to assist in providing context. IDE is invoked to provide useful suggestions, when reaching the pre-defined trigger point. RLPG (Shrivastava et al., 2023) analyses dependencies potentially existing in the target function, then extracts project context that may be conducive to generating the target function. RepoCoder (Zhang et al., 2023b) first slices the source code files in the repository to construct a retrieval library, then uses the generated problem as a query, calculates similarity to obtain project context, and feeds the project context and problem to the model to generate initial results. After obtaining the initial results, the previous generation results are used as queries for further querying to obtain more context. This process is repeated to continuously optimize the generation results. $A^{3}$-CodGen (Liao et al., 2024) first constructs a repository knowledge base covering knowledge of third-party libraries and extracts information on mobile and code context. During the code generation process, LLM first generates an initial version of code based on NL intent, then uses NL intent and the initial version of code to retrieve as much foundational knowledge as possible from the knowledge base for the LLM. Finally, local, global, and third-party library information is integrated into the prompt to allow the model to generate the final result.
However, all the mentioned works treat the latest version of the project repository as the context source, falling into the category of evolving-ignored scenarios. Although in RepoCoder, researchers begin to take notice of this issue by considering all code after the target function in the corresponding file as “future code”, this still falls short of reality and introduces a lack of the context that should be present. Therefore, we propose a novel repository-level code generation benchmark, HumanEvo, an evolving-aware pragmatic benchmark, to better simulate real development scenarios for LLMs.
2.3. LLMs for Code Generation
CodeLlama (Rozière et al., 2024) is further trained based on Llama2 (Touvron et al., 2023). It is capable of understanding and generating code in various programming languages, including but not limited to Python, Java, JavaScript, etc. CodeLlama not only generates complete code but also provides suggestions and optimization solutions for specific code snippets, greatly assisting developers. During the training of the base model, weights are initialized with an equivalent parameter amount of the Llama2 model, followed by training on a dataset of 500 billion tokens of code. StarCoder (Li et al., 2023) is derived from StarCoderBase, which is trained using 10 trillion tokens from The Stack, a vast collection of licensed GitHub repositories with inspection tools and selection exit procedures. The DeepSeek-Coder (Guo et al., 2024) takes into account the often intricate dependency relationships inherent in real-world projects. The model needs to fully consider these dependencies before completing programming tasks. Therefore, to equip DeepSeek-Coder to handle real-world complex projects, the training data is structured to support repository-level comprehension, enhancing the model’s understanding of complex projects. This enables the model to utilize cross-file context effectively. MetaGPT (Hong et al., 2024) constructs a multi-agent collaborative framework, employing Standardized Operating Procedures and involving multiple roles such as programmers and product managers to collaborate on project completion. The results generated by agents are not saved in conversational form but in structured output formats such as charts, files, etc. Communication between agents is also conducted through structured files. MetaGPT also implements an iterative development pipeline, optimizing results through multiple rounds of iteration by recording information generated during the development process by different roles. Similarly, ChatDev (Qian et al., 2023) adopts a multi-agent framework, allocating and assigning different AI agents to various job functions. ChatDev facilitates communication among different role agents, forming a complete software development process where each agent performs its respective tasks while collaborating uniformly. In addition to these models, there are many other excellent models used in the field of code generation (Luo et al., 2024; Christopoulou et al., 2022; Shen et al., 2023; Nijkamp et al., 2023; Chen et al., 2021b; Li et al., 2022; Fried et al., 2023).
2.4. Project Evolution
GitHub-hosted projects undergo continuous evolution over time (Negoita et al., 2019). GitHub provides an open platform where the development progress, issues, and solutions of projects are visible to everyone (Dabbish et al., 2012; Kalliamvakou et al., 2015). When users identify missing features in a project, they typically submit an issue to describe their needs to the developers (Kallis et al., 2021). To address the issue, developers will create a new branch and write code to implement the new feature or fix the bug according to user requirements. Once the code changes are complete, developers commit them to the new branch, push the branch to the GitHub repository, and then initiate a pull request to propose merging these code changes into the main branch (Rahman and Roy, 2014b). Finally, project administrators will review the changes in the pull request and, upon approval, merge them into the main branch.
3. HumanEvo Benchmark
@@
In this section, we provide a detailed introduction to HumanEvo from three perspectives: benchmark overview, benchmark construction pipeline, and benchmark characteristics.
3.1. Benchmark Overview
Figure 2. Task instance overview.
@@ 3.1. Benchmark Overview
HumanEvo consists of a total of 400 task instances, comprising 200 Python programming questions and 200 Java programming questions, respectively. As illustrated in Figure 2, each task instance is associated with a unique identification number. The Owner/Repo field helps identify the GitHub project from which the selected function originates, while the Version field indicates the version of the project during the implementation of this function. Base commit refers to the last commit on the branch before the target function is committed. We use it to restore the project repository to the state before the target function was implemented. The Function declaration of the target function will be provided to the LLM as part of the prompt, along with its corresponding docstring. Since the docstring styles of functions in pragmatic projects vary, with some adopting line-by-line comments and others even lacking corresponding comments, we think it is necessary to rewrite the docstring for each selected function. As depicted in Figure 2, we provide two styles of docstrings. The Brief docstring provides a concise summary of the functionality of the target function, while the Detailed docstring provides a thorough description of the function’s functionality, inputs, and outputs.
3.2. Benchmark Collection Pipeline
Figure 3. HumanEvo construction pipeline.
As depicted in Figure 3, the construction pipeline of HumanEvo consists of six steps, mainly revolving around the collection of high-quality PRs. Here are the reasons why we construct HumanEvo through PRs. We use PRs to construct HumanEvo primarily for the following three reasons: First, since prompting LLMs to generate repository-level functions can be viewed as an incremental development process, and PRs for adding new features exactly mirror the incremental development process in the real world, we choose to have LLMs generate the newly added functions in PRs to simulate real-world development scenarios. Second, the quality of functions in PRs is guaranteed. Before merging the newly added code in PRs into the project, repository administrators conduct a review on it, which allows us to select high-quality task instances directly. Third, the base commit in PRs helps us roll back the project repository to the state before the target function was committed. Finally, by processing PRs appropriately, we can further determine whether the target functions are covered by the project’s testing framework.
3.2.1. Projects selection

### v2

# HumanEvo: An Evolution-aware Benchmark for More Realistic Evaluation of Repository-level Code Generation

@@ Abstract
To evaluate the repository-level code generation capabilities of Large Language Models (LLMs) in complex real-world software development scenarios, many evaluation methods have been developed. These methods typically leverage contextual code from the latest version of a project to assist LLMs in accurately generating the desired function. However, such evaluation methods fail to consider the dynamic evolution of software projects over time, which we refer to as evolution-ignored settings. This in turn results in inaccurate evaluation of LLMs’ performance. In this paper, we conduct an empirical study to deeply understand LLMs’ code generation performance within settings that reflect the evolution nature of software development. To achieve this, we first construct an evolution-aware repository-level code generation dataset, namely HumanEvo, equipped with an automated execution-based evaluation tool. Second, we manually categorize HumanEvo according to dependency levels to more comprehensively analyze the model’s performance in generating functions with different dependency levels. Third, we conduct extensive experiments on HumanEvo with seven representative and diverse LLMs to verify the effectiveness of the proposed benchmark. We obtain several important findings through our experimental study. For example, we find that previous evolution-ignored evaluation methods result in inflated performance of LLMs, with performance overestimations ranging from 10.0% to 61.1% under different context acquisition methods, compared to the evolution-aware evaluation approach. Based on the findings, we give actionable suggestions for more realistic evaluation of LLMs on code generation. We also build a shared evolution-aware code generation toolbox to facilitate future research. The replication package including source code and datasets is available at https://github.com/DeepSoftwareAnalytics/HumanEvo.
I Introduction
@@
In recent years, the LLM-based code generation task has drawn widespread attention [36, 37, 38, 52, 53, 54, 24, 25, 75, 76]. Many code LLMs [8, 9, 10, 39, 40, 51, 29, 30, 31, 17, 32, 33, 18] are being employed to empower programming assistants, which are playing a vital role in practical software development [70, 71, 72, 73]. Recently, several repository-level code generation benchmarks such as CoderEval [13], RepoBench [14], EvoCodeBench [64], and RepoEval [6] have emerged to better simulate real-world development scenarios. They sample functions from real-world projects as programming tasks, aiming to reflect the performance of LLMs in actual development by prompting them to generate these functions according to the target function description and project context.
Despite the fact that real-world projects are inherently dynamic and evolve over time [55, 56, 57], current code generation benchmarks often overlook this critical aspect. They tend to treat the latest version of the repository as context source for evaluation, a situation we term as the evolution-ignored phenomenon. This evaluation approach fails to accurately capture true nature of real software development scenarios where projects continuously evolve over time, which means that the project context that developers face when writing different code changes constantly. Therefore, to accurately assess the repository-level coding capabilities of LLMs, it is imperative to factor in this evolution aspect. Specifically, when prompting LLMs to generate different target functions, the project context provided should mirror the one available to developers at the time of the target function’s creation. Unfortunately, evolution-ignored evaluation in previous benchmarks provides LLMs with an inaccurate and potentially misleading programming environment, leading to serious issues. We have found concrete instances of issues that stem from the evolution-ignored setting in real GitHub projects, which are detailed in Section II.
@@
Benchmark HumanEvo. To fill this gap, we construct HumanEvo, a novel evolution-aware repository-level code generation benchmark, which better simulates real-world development processes and reflects the evolution nature of projects over time. The evaluation process of HumanEvo is as follows: to ensure that the context we provide to LLMs mirrors the context available to the programmer when writing the target function (neither including additional code introduced by subsequent version changes nor omitting code that may have been updated or deleted due to version change), we roll back the entire repository to the state before the target code was committed and evaluate the code generation performance of LLMs at function granularity on the rolled-back repository.
@@
We conduct a rigorous data construction pipeline to construct HumanEvo. We start with selecting high-quality projects and then collecting a large number of pull requests (PRs) from these projects. Then, we perform an initial filtering on the crawled PRs to meet three attributes to guarantee the quality of the newly added functions in PRs. Additionally, to ensure the reliability of the HumanEvo’s test suite, we establish a runtime environment for each PR and execute the corresponding project’s testing framework to verify that the newly added functions are covered by the test suite. Finally, we obtain 400 task instances, 200 for Python and 200 for Java. For each task instance in HumanEvo, we record its corresponding PR’s metadata to roll back the repository to the state prior to the target function’s commit.
Empirical Study. Based on the new benchmark HumanEvo, we perform the first study to reveal LLMs’ repository-level code generation capability in the evolution-aware setting. In particular, we conduct an extensive evaluation of 7 mainstream LLMs, including both open-source ones (CodeLlama-7B, CodeLlama-13B, CodeLlama-34B [8], DeepSeekCoder-6.7B, DeepSeekCoder-33B [10]) and closed-source ones (GPT-4 [16] and GPT-3.5-Turbo [74]) on HumanEvo.
According to the empirical results, we have the following findings: ① We find that all the studied LLMs show much worse performance on our evolution-aware setting. Previous evolution-ignored evaluation would lead to the inflated performance of LLMs, ranging from 10.0% to 61.1%, revealing how previous evaluation methods provided the models with a deviated programming scenario from reality. ② We find that the impact of the evolution-ignoring setting is more severe for target code with more complex dependencies. Specifically, compared to the evolution-ignored setting, the success rate of LLMs in generating functions with intra-class dependencies under the evolution-aware setting has been observed to decrease by approximately 14.0% on average, while for functions with inter-class dependencies, the decline is 30.9% on average. ③ We investigate how the performance of LLMs change as the repository evolve. Experimental results show that in the evolution-ignored setting, in general, models’ performance gradually deviates from the real performance as the project evolves. ④ We find that , under the evolution-ignored setting, utilizing docstrings in different styles as input consistently leads to inflated performance. Experimental results indicate that LLMs’ code generation performance improves when given a detailed docstring compared to a brief one.
This work makes the following key contributions:
@@
  • We have identified a common flaw in previous repository-level code generation benchmarks: they are evolution-ignored, causing inaccurate simulations of real-world development scenarios.
@@
  • We introduce a new benchmark HumanEvo, an evolution-aware repository-level code generation benchmark that addresses the evolution-ignored issue, to conduct a comprehensive empirical study on the performance of LLMs in repository-level code generation.
@@
  • Through extensive experimentation, we substantiate the evolution-ignored situation leads to inflated performance of the models, with LLMs showing a performance inflation ranging from 10.0% to 61.1%.
@@
  • For convenient usage of HumanEvo, we release all data, code, and the evaluation platform with Docker images that provide runtime environments for all projects, facilitating automated execution of all test suites. The replication package is provided at https://anonymous.4open.science/r/HumanEvo/.
Figure 1: Example 1: future context leakage.
Figure 2: Example 2: useful context missing.
II Motivating Examples
In this section, we firstly elaborate on how real repository evolves over time. Then, we present two motivating examples [^1] The motivating examples are from the real-world project https://github.com/django/django that showcase the specific issues that can emerge under the evolution-ignored setting.
In real-world development, projects undergo continuous evolution over time [60, 61, 62]. During the development phase of a project, programmers need to develop the project (i.e., by adding new features, removing existing code, adding test code, etc) based on the project requirements. In the maintenance phase, programmers continuously monitor the project and promptly fix any bugs that arise. These activities result in varying degrees of modifications to the project, causing the project to evolve over time [59, 63]. For a realistic and accurate evaluation of LLMs’ code generation capabilities, this evolution nature should be considered. Otherwise, severe issues under the evolution-ignored setting would emerge. We have identified concrete instances of issues, which we refer to as future context leakage and use context missing, which we will illustrate with the following two examples.
Example 1. Figure 1 illustrates the future context leakage issue that can arise from the evolution-ignored setting. This example involves a target function that was originally committed on Jan 23, 2019. In the canonical RAG code generation pipeline, a retriever finds similar code from the corresponding repository (which is the latest version in the evolution-ignored setting) to help LLMs generate the target code. However, the retriever yields a function committed on Sep 2, 2022, a date that is later than the target function. This process is illogical because it is implausible for a developer to access code that was not yet written at the time of the creation of the target function. Therefore, providing this code snippet as context to the model is a future context leak. While in this example the LLM generates the correct code, the reliability is questionable as it is uncertain whether it can still generate correctly without the help of this future code.
Example 2. Figure 2 illustrates the useful context missing issue that may arise in the evolution-ignored setting. In this example, the function committed on Jan 7, 2015 is the target function. In the same RAG code generation pipeline as in Example 1, the retrieved code snippet is not particularly useful and LLM fails to generate the correct code. However, when the developer wrote that code, there was actually a piece of code in the repository that was highly similar to the target code and could have been retrieved as context, which would have likely helped generate the correct code. But this code snippet was deleted during the evolution process. As a result, under the evolution-ignored setting, this piece of code did not exist in repository context. While in this example the LLM generates wrong code, it is possible that providing LLM with this useful context might have enhanced its performance.
III HumanEvo Benchmark
@@
In this section, we introduce our new benchmark HumanEvo in detail. We present the benchmark overview (Section III-A), benchmark construction pipeline (Section III-B), and benchmark characteristics (Section III-C).
III-A Benchmark Overview
Figure 3: Task instance overview.
@@ III-A Benchmark Overview
HumanEvo consists of a total of 400 task instances, comprising 200 Python programming tasks and 200 Java programming tasks, respectively. Figure 3 shows an example of a task instance overview. Each task instance is associated with a unique identification number (Instance id). The Owner/Repo field helps identify the GitHub repository from which the selected function originates, while the Version field indicates the version of the repository according to the implementation time of this function. Base commit refers to the last commit on the branch before the target function was committed. We use it to restore the repository to the state before the target function was implemented. The Function declaration of the target function will be provided to the LLM as part of the prompt, along with its corresponding docstring. Since the docstring styles of functions in pragmatic projects vary, with some adopting line-by-line comments and others even lacking corresponding comments, we think it is necessary to rewrite the docstring for each selected function. As depicted in Figure 3, we provide two styles of docstrings. The Brief docstring provides a concise summary of the functionality of the target function, while the Detailed docstring provides a thorough description of the function’s functionality, inputs, and outputs.
III-B Benchmark Construction Pipeline
Figure 4: HumanEvo construction pipeline.
As depicted in Figure 4, the construction pipeline of HumanEvo consists of six steps, namely, (1) Project Selection; (2) Pull Request Crawling; (3) Attribute-based Filtering; (4) Execution-based Filtering; (5) Function Selection; and (6) Human Annotation.